OpenStudy (anonymous):

What is the measure of one interior angle & exterior angle in a regular heptagon? please explain i dont get it

OpenStudy (anonymous):

The sum of all exterior angles of a polygon is 360 degrees. A heptagon has 7 sides. Since it is a regular heptagon, all sides and angles are equal. so each exterior angle will have 360/7 = 51.42857143 degrees.

OpenStudy (anonymous):

As far as each interior angle of a REGULAR heptagon, it is 180(n-2)/n so it is 180(5)/7 = 128.5714286 degrees

OpenStudy (tkhunny):

You could notice that interior and exterior angles form linear pairs. 180 - 51.4285714 = 128.5714286